### v1

# The Giroux Correspondence in dimension 3

Abstract.
In [LV] the authors proved the Giroux Correspondence for tight contact $3$-manifolds via convex Heegaard surfaces. Simultaneously, [HBH] gave an all-dimensions proof of the Giroux Correspondence by generalising convex surface theory to higher dimensions. This paper uses a key result from [HBH] and also [Tia] to complete the $3$-dimensional proof for arbitrary (not necessarily tight) contact 3-manifolds. This presentation features low-dimensional techniques and further clarifies the relationship between contact manifolds and their Heegaard splittings that was introduced in [LV].
1. Introduction
Topology and contact geometry enjoy a remarkable relationship in dimension $3$. Contact structures are defined in the smooth category with differential geometric language, and yet they mirror fundamental topological properties of the underlying $3$-manifold. An open book decomposition is arguably the most striking incarnation of this relationship: up to diffeomorphism, a contact manifold is determined simply by a surface mapping class. For the last twenty years, 3-dimensional contact geometry has been shaped by the equivalence between contact structures and open books first proposed by Giroux in the early 2000s:
Theorem 1.1.
[Giroux Correspondence [Gir02]] Two open book decompositions support isotopic contact structures if and only if they admit a common positive stabilisation.
Because topology and contact geometry are so closely entwined, it’s natural to probe whether topological techniques can be squeezed to produce meaningful contact geometric outputs. In this spirit, in 2023 the authors developed new tools to study $3$-dimensional contact manifolds via their Heegaard decompositions.
We used the enhanced relationship between contact structures and Heegaard splittings to prove the Giroux Correspondence in the case of tight contact $3$-manifolds [LV]. Simultaneously, Breen-Honda-Huang used generalised convex surface theory to prove the Giroux Correspondence for all contact structures and in all dimensions [HBH]. While the work of Breen-Honda-Huang fully covers the $3$-dimensional case, the generality of their techniques lends their proof a different flavour. This paper completes the 3-dimensional story, using an essential result of [HBH] and also [Tia] to bridge the gap to overtwisted manifolds that limited the scope of our previous paper. We believe this work will make the proof more accessible to a low-dimensional audience; in particular, it offers a useful illustration of convex surface techniques in concert with Heegaard splittings.
1.1. Heegaard splittings and open book decompositions
Any open book decomposition of a manifold gives rise to a canonical Heegaard decomposition. Heegaard splittings produced thus are called convex Heegaard splittings, and one may directly recover an open book supporting $\xi$ from any convex Heegaard splitting of $(M,\xi)$. The innovation in [LV] is a procedure called refinement that produces an open book from a more general class of Heegaard splittings of a contact manifold. Although this open book decomposition is not uniquely defined, any two open books produced thus admit a common positive stabilisation – exactly the equivalence appearing in the Giroux Correspondence. It follows that the Giroux Correspondence can be rephrased as a statement about convex Heegaard splittings (Theorem 3.3). Our proof of the Giroux Correspondence for tight contact manifolds relied on showing that many operations one might naturally perform on a Heegaard splitting of a contact manifold in fact preserve the positive stabilisation class of the associated open book.
However, these techniques were insufficient in the case of overtwisted manifolds. In order to assign an open book decomposition to a contact manifold with a Heegaard splitting, our approach required both handlebodies to be tight. When two convex Heegaard surfaces are isotopic, there is a sequence of convex Heegaard surfaces interpolating between them. However, even when both original Heegaard splittings have tight handlebodies, intermediate Heegaard splittings in the sequence may nevertheless have overtwisted handlebodies. The present work bypasses this stumbling block by considering highly stabilised Heegaard splittings where the handlebodies are always tight. We introduce the term “bridging” for a particular method of constructing such a splitting, and we show that different choices made in constructing a bridge splitting preserve the positive stabilisation class of the associated open book. This allows us to prove that any pair of convex Heegaard splittings for a fixed contact manifold correspond to open books that admit a common positive stabilisation.
1.2. A note on notation
The arguments in this paper involve sequences of Heegaard splittings of varying types, each of which is denoted by some sort of “$\mathcal{H}$”. We have attempted to define notation locally and maintain these choices consistently, but we briefly outline the system here for reference. Every Heegaard surface $\Sigma$ is required to be convex; Heegaard splittings subject to no further restrictions are denoted by $\mathcal{H}=(\Sigma,U,V)$, where $U$ and $V$ are the handlebodies, oriented so that $\partial U=\Sigma$. Elements of a pair of Heegaard splittings are distinguished by the addition of a prime: $\mathcal{H}=(\Sigma,U,V)$ and $\mathcal{H}^{\prime}=(\Sigma^{\prime},U^{\prime},V^{\prime})$. Following the notation in [LV], the refinement of $\mathcal{H}$ is denoted by ~$\widetilde{\mathcal{H}}=(\widetilde{\Sigma},\widetilde{U},\widetilde{V})$.
Starting in Section 3.2, we will stabilise Heegaard splittings by attaching certain contact $1$-handles found in the complement of fixed handlebodies. We will call this process bridging, and we denote the bridge of $\mathcal{H}$ by ^$\widehat{\mathcal{H}}=(\widehat{\Sigma},\widehat{U},\widehat{V})$. When we bridge only a single handlebody of the original splitting, we denote the result by ^$\widehat{\mathcal{H}}^{U}$ or ^$\widehat{\mathcal{H}}^{V}$, as appropriate. Distinct bridged splittings may be distinguished by primes: $\widehat{\mathcal{H}}$ and $\widehat{\mathcal{H}}^{\prime}$.
Acknowledgment
This research was supported in part by the Austrian Science Fund (FWF) P 34318. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. The authors would also like to thank Ko Honda and Matthias Scharitzer for helpful discussions on the topic.
2. Background
We assume the reader is familiar with basic notions in contact geometry, but we use this section to collect some tools and terminology that will be essential for the rest of the paper. Section 2.1 provides a quick review of convex surface theory and introduces weak contactomorphism (Definition 2.1) as an equivalence relation on contact manifolds with convex boundary. Section 2.2 describes the relationship between open book decompositions and convex Heegaard splittings (Definition 2.4) of a contact manifold. Section 2.3 presents the contact handle model for bypass attachment. Finally, Section 2.4 states the key theorem borrowed from [HBH] and [Tia] which classifies distinct bypass decompositions of a product cobordism. Also, Lemma 2.8 gives a new perspective on the familiar operation of bypass rotation.
2.1. Convex surfaces
The reader looking for a more thorough introduction to convex surface theory is referred to [Mas14] or the lecture notes [Etn04, Hon].
Recall that a vector field $\mathcal{V}$ in a contact manifold $(M,\xi)$ is contact if the flow of the vector field preserves $\xi$. A surface is convex when there exists a contact vector field $\mathcal{V}$ transverse to $\Sigma$. The essential feature of a convex surface $\Sigma$ is that it provides a combinatorial characterisation of the contact structure in an $I$-invariant neighbourhood $\nu(\Sigma)$; the product structure on $\nu(\Sigma)$ is provided by the flow of $\mathcal{V}$. The dividing set on $\Sigma$ is a separating multicurve $\Gamma_{\Sigma}$ consisting of the points where $\mathcal{V}$ lies in $\xi$. Changing the vector field that certifies $\Sigma$ as convex changes $\Gamma_{\Sigma}$ only by isotopy, and any isotopy keeping $\Sigma$ convex also preserves $\Gamma_{\Sigma}$ up to isotopy. Furthermore, any $\Gamma^{\prime}$ isotopic to $\Gamma_{\Sigma}$ may be realised as a dividing set after performing an isotopy of $\Sigma$ within an $I$-invariant neighbourhood.
Convex surface theory is particularly useful for studying contact manifolds with boundary, and we will assume throughout that the boundary of $(M,\xi)$ is convex. Since the dividing set characterises $\xi$ near $\partial M$, we would like to be able to glue $(M,\xi,\Gamma_{\partial M})$ to $(M^{\prime},\xi^{\prime},\Gamma_{\partial M^{\prime}})$ whenever there is a diffeomorphism taking $(\partial M,\Gamma_{\partial M})$ to $(\partial M^{\prime},\Gamma_{\partial M^{\prime}})$. Because we are working in the smooth category, gluing is achieved by identifying contact neighbourhoods of the boundaries; for later convenience, we introduce the following equivalence relation on contact manifolds with boundary:
Definition 2.1.
[LV, Definition 2.3.] Two contact manifolds $(M,\xi)$ and $(M^{\prime},\xi^{\prime})$ with convex boundary are weakly contactomorphic if there is a contact embedding $\iota\colon M\hookrightarrow M^{\prime}$ such that $\iota(\partial M)$ is convexly isotopic to $\partial M^{\prime}$.
Remark 2.2.
We note that if $\iota:M\rightarrow M^{\prime}$ is a contact embeddeding as above, then $M\setminus\iota({M^{\prime}})$ is a weak identity morphism in the sense of [Tia].
Henceforth, we will be interested in contact manifolds with boundary only up to weak contactomorphism. Two (weak contactomorphism classes of) manifolds may be glued along diffeomorphic boundary components if and only if there is a diffeomorphism respecting the dividing sets. Gluing is well defined as an operation on weak contactomorphism classes of manifolds; see [LV, Proposition 2.8.].
As a special case of weakly contactomorphic manifolds, we also introduce the following:
Definition 2.3.
[LV, Definition 2.4.] Two embedded codimension-00 submanifolds $N_{1},N_{2}\subset(M,\xi)$ with convex boundary are weakly (contact) isotopic if there is an isotopy $N_{s}$ between them such that $\partial N_{s}$ remains convex throughout.
2.2. Open book decompositions of contact manifolds
Let $(B,\pi)$ denote an open book decomposition of $M$. That is, $B$ is an oriented link and $\pi:M\setminus B\rightarrow S^{1}$ is a fibration of the complement of $B$ by Seifert surfaces for $B$. When $M$ is equipped with a contact structure $\xi$, we assume any open book decomposition supports $\xi$ in the sense of [TW75].
An open book decomposition $(B,\pi)$ induces a Heegaard splitting $(\Sigma,U,V)$ where each handlebody consists of half the pages: $U=\overline{\pi^{-1}[0,1/2]}$ and $V=\overline{\pi^{-1}[1/2,1]}$. The Heegaard surface
$\Sigma=\overline{\pi^{-1}(0)}\cup-\overline{\pi^{-1}(1/2)}$
is naturally convex with dividing curve $\Gamma_{\Sigma}=B=\partial\big{(}\pi^{-1}(0)\big{)}$. An arc properly embedded on a page of the open book produces a properly embedded disc in the corresponding handlebody, and in fact, such discs characterise the Heegaard splittings induced by open books.
More precisely, a disc $A$ properly embedded in a contact manifold with convex boundary $(\Sigma,\Gamma_{\Sigma})$ is a product disc if $|\partial A\cap\Gamma_{\Sigma}|=2$. A disc system for a handlebody is a set of properly embedded discs that cut it into a ball. Torisu introduced the notion of a convex Heegaard splitting [Tor00]:
Definition 2.4.
A Heegaard splitting $(\Sigma,U,V)$ of $(M,\xi)$ is convex if
  • $\Sigma$ is convex with dividing curve $\Gamma_{\Sigma}$;
  • $\xi|_{U}$ and $\xi|_{V}$ are both tight; and
  • each of the manifolds $U$ and $V$ admits a system of product discs.
The discs in a convex Heegaard splitting determine a product structure on each handlebody, and hence, an open book decomposition with binding $\Gamma_{\Sigma}$. Moreover, this relation is one-to-one up to isotopy.
Proposition 2.5.
[LV, Proposition 3.4.] Let $(M,\xi)$ be a contact manifold and let $(B,\pi)$ and $(B^{\prime},\pi^{\prime})$ be open books supporting $\xi$. Then the induced convex Heegaard splittings are isotopic via an isotopy keeping the Heegaard surface convex if and only if $(B,\pi)$ and $(B^{\prime},\pi^{\prime})$ are isotopic through a path of open books supporting $\xi$.
Because a convex Heegaard splitting determines an open book decomposition, a convex splitting allows one to reconstruct the contact structure. Specifically, suppose $(\Sigma,\Gamma_{\Sigma},\{\alpha_{i}\},\{\beta_{i}\})$ is a convex Heegaard diagram for $(M,\xi)$ where the meridional circles bound product discs in the respective handlebodies and the restriction of $\xi$ to each handlebody is known to be tight. Then up to contact isotopy, there is a unique contact structure structure compatible with this data.
2.3. Bypass attachment
Next, we turn to an operation that changes the weak contactomorphism class of a contact manifold while preserving the topological type. Adopting the perspective of [HKM09, Ozb11], we define bypass attachment as an operation performed by gluing contact handles.
As in the topological setting, a contact $k$-handle is a $D^{k}\times D^{3-k}$ glued along $\partial D^{k}\times D^{3-k}$ to the boundary of an existing manifold $M$. Additionally, each $k$-handle is required to smooth to a tight $3$-ball, and the dividing set on the attaching region is prescribed. A $1$-handle attachment is specified by an $S^{0}$ on the dividing set $\Gamma_{\partial M}$, while a $2$-handle attachment is specified by an $S^{1}$ that intersects $\Gamma_{\partial M}$ twice transversely. Local models for handle attachment are well established in the literature; see, for example, [Gir91] or [Ozb11]. Although each handle is a cornered manifold, the gluing process is defined so that the manifold is smooth after each attachment.
Let $(\Sigma,\Gamma_{\Sigma})$ be a convex surface. A Legendrian arc $c$ embedded in $\Sigma$ is a bypass arc if $\partial c$ lies on $\Gamma_{\Sigma}$ and the interior of $c$ intersects $\Gamma_{\Sigma}$ once transversely. As explained next, we may associate a new contact manifold with convex boundary to any bypass arc on the convex boundary of a contact manifold.
Figure 1. Left: a bypass half-disc attached along the arc $c$. Centre: the attached $1$-handle is a regular neighbourhood of an arc Legendrian isotopic to $c$. Right: attaching the shaded $2$-handle along $c\cup c^{\prime}$ completes the bypass attachment.
[…]
Figure 2. Left: Local model for a bypass attachment arc. Centre: The new dividing set after attaching a bypass along the arc shown on the left. Right: A trivial bypass arc. Attaching a bypass along the arc shown preserves the dividing set up to isotopy.
Suppose that $(M,\xi,\Gamma_{\partial M})$ is a contact manifold with convex boundary and let $c$ be a bypass arc on $\partial M$. Then the weak contactomorphism class of $M\cup(h^{1}\cup h^{2})$ is preserved under isotopies of $c$ through bypass arcs.
A bypass is trivial when the bypass arc $c$ on $\Sigma$ matches the one shown on the right in Figure 2. That is, the arc $c$ turns left after the interior crossing with $\Gamma_{\Sigma}$ to cobound a bigon with the same arc of $\Gamma_{\Sigma}$.
Attaching a trivial bypass preserves the convex isotopy class of the surface. If $\Sigma$ bounds a contact submanifold $(N,\xi)$, then attaching a trivial bypass also preserves the weak contact isotopy class of $N$. As shown in [HKM07, Lemma 5.1.], when the bypass arc $c$ is trivial on $\Sigma\subset(M,\xi)$, the corresponding bypass half-disc $D$ always exists as a submanifold of $M$.
2.4. Bypass decompositions
Bypasses allow us to classify distinct contact structures on a fixed topological manifold. Adopting cobordism notation, we write $\partial_{+}M$ (respectively, $\partial_{-}M$) to distinguish the outward-oriented (respectively, inward-oriented) boundary of a topological $M$.
Theorem 2.6.
[Hon02, Section 3.2.3] Let $(\Sigma\times I,\xi)$ be a topological product with convex boundary. Then there exists a sequence of bypasses $X_{i}:=h_{i}^{1}\cup h_{i}^{2}$ such that $(\Sigma\times I,\xi)$ is weakly contactomorphic to
$\nu(\Sigma)\cup\bigcup_{i}X_{i},$
where $\nu(\Sigma)$ is an $I$-invariant half-neighbourhood of $\Sigma\times\{-1\}$ and the bypass arc $c_{j}$ for $X_{j}=h_{j}^{1}\cup h_{j}^{2}$ lies on
$\partial_{+}\big{(}\nu(\Sigma)\cup\bigcup_{i=1}^{j-1}X_{i}\big{)}.$
Theorem 2.6 has an alternative presentation as a decomposition theorem. Given a contact manifold $(\Sigma\times I,\xi)$ with convex boundary, consider an embedding
$\iota:\nu(\Sigma)\cup\bigcup_{i}X_{i}\hookrightarrow(\Sigma\times I,\xi)$
as in Definition 2.1. It follows that contact handles $h_{i}^{1}$ and $h_{i}^{2}$ may be identified as submanifolds of the original $(\Sigma\times I,\xi)$. This identifies a sequence of bypass arcs $c_{j}$ on the embedded convex surfaces $\Sigma_{j-1}:=\iota(\partial_{+}\big{(}\nu(\Sigma)\cup\bigcup_{i=1}^{j-1}X_{i}\big{)}\big{)}$. Similarly, there exist embedded bypass half-discs $D_{i}$ in $\Sigma\times I$. The surface $\Sigma_{j}$ is constructed by isotoping $\Sigma_{j-1}$ across a neighbourhood of $D_{j}$.
We write $(\Sigma\times I,\xi)=X_{1}\ast X_{2}\ast\dots\ast X_{f}$ to denote this decomposition of $(\Sigma\times I,\xi)$.
The factorisation provided by this theorem is not unique. As mentioned before, one can isotope any bypass attaching arc through bypass arcs. In addition, there are two further moves that change the decomposition while preserving the contact manifold: Trivial Insertion and Far Commutation.
[TI]: Trivial Insertion Given a bypass decomposition $X_{1}\ast\dots\ast X_{f}$, let $c_{T}$ denote a trivial bypass arc on $\Sigma_{j}$. Then a bypass attachment along $c_{T}$ may be inserted into the decomposition:
[…]
[FC]: Far Commutation Suppose that the bypass arcs $c_{j}$ and $c_{j+1}$ for $X_{j}$ and $X_{j+1}$ are disjoint. Then in any decomposition in which these correspond to consecutive bypass attachments, the order in which these are performed may be exchanged:
[…]
The main theorem of both [Tia] and [HBH] establishes that Trival Insertion and Far Commutation suffice to relate any two bypass decompositions of topologically trivial products:
Theorem 2.7.
[HBH, Theorem 3.1.2 ] Any two bypass decompositions of $(\Sigma\times I,\xi,\Gamma_{\Sigma})$ are related by isotopy of the bypass arcs and a finite iteration of Trivial Insertion and Far Commutation moves.
The dimension 3 case was the Main Theorem in [Tia], although presented in different language.
The term bypass rotation is used in several different, but related, senses in the literature. Here, we consider the phenomenon as presented in [HKM07] and recall the proof that it is a consequence of the above two moves. Let $c$ be a bypass arc on $\Sigma$. Let $c_{T}$ be another bypass arc as in Figure 3; i.e., subsegments of $c$ and $c_{T}$ together with two segments of $\Gamma_{\Sigma}$ cobound a rectangle $R$ that is to the right of $c$ when $c$ is oriented as the boundary of $R$. Let $R^{\prime}\subset\Sigma$ be a neighbourhood of $R\cup c\cup c_{T}$ shown shaded in the right hand picture of Figure 3. The arc $c_{T}$ is said to be a rotation of the bypass arc $c$.
Figure 3. Left: The bypass arc $c_{T}$ is a rotation of the bypass arc $c$. Left centre: After attaching a bypass along $c$, $c_{T}$ becomes a trivial bypass arc. Right centre: The dividing set is preserved up to isotopy by attaching a bypass to $c_{T}$. Right: The bypass disc attached along $c_{T}$ can be constructed in a neighbourhood of $R^{\prime}$ and the bypass disc for $c$.
Lemma 2.8.
Suppose that a bypass disc $D$ is attached to the arc $c$ on some surface $\Sigma\subset(M,\xi)$, as shown in Figure 3. Then there also exists a bypass disc $D_{T}$ attached to $\Sigma$ along $c_{T}$. Furthermore, $D_{T}$ is contained in a neighbourhood of $D$ and $R^{\prime}$.
The second claim of the lemma will be useful in the proof of Proposition 3.8 below.
Proof.
The existence claim follows from a formal convex surface theory argument. After a bypass is attached along $c$, $c_{T}$ is a trivial bypass arc on the new surface. Since trivial bypasses always exist, it follows that there exists a bypass disc $D_{T}$ attached along $c_{T}$. The [TI] move then implies that $\nu(\Sigma)\cup X_{c}$ can be factored as $\nu(\Sigma)\cup(X_{c}\ast X_{c_{T}})$. However, the bypass arcs $c$ and $c_{T}$ are disjoint on $\Sigma$, so the associated bypasses commute. Applying the [FC] move gives the factorisation $\nu(\Sigma)\cup(X_{c_{T}}\ast X_{c})$. It follows that the bypass along $c_{T}$ can be directly attached onto $\Sigma$.
The next argument is illustrated in Figure 4.
Figure 4. Attach a $1$-handle $h_{T}^{1}$ along a push-off of the Legendrian curve $d$. Then isotope the disc cobounded by $h_{T}^{1}$ and $d$ across the product disc $A$ to produce a bypass disc $D$ cobounded by $h_{T}^{1}$ and $c_{T}$.
To construct $D_{T}$ directly, consider a Legendrian realisation of the arc labeled $d$ in Figure 4. Perform a Legendrian isotopy so that $d$ is properly embedded in $\nu(\Sigma)\cup X_{c}$. Since $d\cap\Gamma_{\Sigma}=\partial d$, a regular neighbourhood of this push-off is a contact $1$-handle $h_{T}^{1}$ shown in bold on the first figure. Observe that $h_{T}^{1}$ is attached to $\Sigma$ along $\partial c_{T}$, but the blue disc it cobounds with $d$ is not a bypass disc. However, we may isotope this disc across the product disc $A$ shown in the centre figure to get a bypass disc $D_{T}$ cobounded by $h_{T}^{1}$ and $c_{T}$. ∎
3. Heegaard Splittings and the Giroux Correspondence
With useful tools and vocabulary established, we now turn to proving the Giroux Correspondence. As a first step, Section 3.1 summarises results from [LV] that allow us to restate the theorem about a pair of open books as a theorem about a pair of Heegaard splittings. Section 3.2 introduces bridging as an operation to stabilise a Heegaard splitting of a contact manifold. After establishing some properties of bridge splittings, we relate them to refinements in Section 3.3 to prove the main result.
3.1. Tight Heegaard splittings and refinement
As described in Section 2.2, each convex Heegaard splitting of a contact manifold corresponds to an open book decomposition supporting the contact structure. Here we recall a key idea from [LV] that extends the class of Heegaard splittings of a contact manifold which determine open books.
Definition 3.1.
The Heegaard splitting $(\Sigma,U,V)$ of $(M,\xi)$ is tight if $\Sigma$ is convex and the restriction of $\xi$ to each of $U$ and $V$ is tight.
The definition of a convex Heegaard splitting requires a system of meridional discs each intersecting $\Gamma_{\Sigma}$ twice. In contrast, meridional discs in a tight Heegaard splitting may intersect $\Gamma_{\Sigma}$ an arbitrary number of times. However, a process called refinement allows us to construct a convex splitting from a tight one.
Refinement was introduced in [LV, Section 5.] . Roughly speaking, refinement stabilises the tight Heegaard splitting by drilling along Legendrian curves embedded in meridional discs. The drilling process cuts each non-product disc into subdiscs, and the curves are chosen so that these subdiscs are product discs for the stabilised Heegaard splitting $\widetilde{\mathcal{H}}$. See Figure 5. This refinement $\widetilde{\mathcal{H}}$ is necessarily convex, so via Lemma 2.5 one may view refinement as a process which associates an open book decomposition to a tight Heegaard splitting of $(M,\xi)$.
Figure 5. To refine a tight Heegaard splitting $\mathcal{H}$, stabilise it via tunnels drilled through non-product discs to produce a convex splitting $\widetilde{H}$. Here, the meriodonal disc $A$ with $|\partial A\cap\Gamma_{\Sigma}|=4$ in $\mathcal{H}$ splits into a pair of product discs $A_{1}$, $A_{2}$ in $\widetilde{\mathcal{H}}$.
There are choices involved in constructing a refinement (e.g., discs, tunnels), but these preserve the positive stabilisation class of the associated open book.
Theorem 3.2.
[LV, Lemma 3.8 and Theorem 5.10] Suppose that $\widetilde{\mathcal{H}}$ and $\widetilde{\mathcal{H}}^{\prime}$ are refinements of the tight Heegaard splitting $\mathcal{H}$. If $(B,\pi)$ and $(B^{\prime},\pi^{\prime})$ are the open book decompositions associated to $\widetilde{\mathcal{H}}$ and $\widetilde{\mathcal{H}}^{\prime}$, respectively, then $(B,\pi)$ and $(B^{\prime},\pi^{\prime})$ admit a common positive stabilisation.
In light of Theorem 3.2, we define a positive stabilisation of a tight Heegaard splitting as a topological stabilisation of the Heegaard splitting which induces a positive open book stabilisation on the open books associated to the refinements.
This new language permits us to restate the hard direction of the Giroux Correspondence:
Theorem 3.3 (3-dimensional Giroux Correspondence).
Any two convex Heegaard splittings of $(M,\xi)$ admit a common positive stabilisation.
Much of the work of [LV] was devoted to showing that various natural operations one may perform on a tight Heegaard splitting are in fact (sequences of) positive stabilisations.
Positive stabilisations may also be characterised directly. Choose an arc properly embedded in one of the handlebodies that is Legendrian isotopic to an arc properly embedded in the closure of $\Sigma\setminus\Gamma_{\Sigma}$. Adding a standard contact neighborhood of this arc to the other handlebody is a positive stabilisation of the original Heegaard splitting. In fact, a related operation has already appeared in the proof of Lemma 2.8. Suppose that in Lemma 2.8 the surface $\Sigma$ is taken to be a Heegaard surface for a tight Heegaard splitting. In this case, the $1$-handle $h^{1}_{T}$ is a neighbourhood of a Legendrian push-off of $d$, so adding $h_{T}^{1}$ to the handlebody bounded by $\Sigma$ is a positive stabilisation of the Heegaard splitting.
The proof of Lemma 3.10 in [LV] shows that in a convex Heegaard splitting, attaching any bypass $1$-handle is a positive stabilisation. The following more general statement will be used later.
Theorem 3.4.
[LV, Lemma 7.1] Let $\mathcal{H}=(\Sigma,U,V)$ and $\mathcal{H}^{\prime}=(\Sigma^{\prime},U^{\prime},V^{\prime})$ be tight Heegaard splittings of $(M,\xi)$ and assume that $U^{\prime}$ is obtained from $U$ by attaching a single bypass $h^{1}\cup h^{2}$. Let
′′$\mathcal{H}^{\prime\prime}=\big{(}\partial(U\cup h^{1}),U\cup h^{1},M\setminus\text{int}(U\cup h^{1})\big{)}$
be the Heegaard splitting formed by attaching only the $1$-handle associated to this bypass. Then the refinements of $\mathcal{H}$, $\mathcal{H}^{\prime}$ and $\mathcal{H}^{\prime\prime}$ all admit a common positive stabilisations.
Remark 3.5.
Refining a convex Heegaard splitting with respect to a set of product discs preserves the splitting. However, a convex splitting $\mathcal{H}$ will in general also have non-product meridional discs. Refining $\mathcal{H}$ with respect to some set of non-product discs will produce a new convex splitting $\widehat{\mathcal{H}}$; then $\mathcal{H}$ and $\widehat{\mathcal{H}}$ admit a common positive stabilisation.
3.2. Bridge Heegaard splittings
Given a tight Heegaard splitting $\mathcal{H}$, the refinement process described above stabilises $\mathcal{H}$ to produce a convex Heegaard splitting $\widetilde{\mathcal{H}}$. In this section we introduce an alternative method to turn any smooth Heegaard splitting $\mathcal{H}$ into a convex Heegaard splitting $\widehat{\mathcal{H}}$. The new Heegaard splitting $\widehat{\mathcal{H}}$ is called the bridge of $\mathcal{H}$.
A bouquet of circles is a skeleton of a handlebody if the handlebody deformation retracts onto the bouquet. When the handlebody is equipped with a contact structure, we may additionally require the bouquet to be Legendrian. Throughout, we will let $K_{U}$ and $K_{V}$ denote Legendrian skeletons of $U$ and $V$, respectively, and we let $\nu(K_{U})$ and $\nu(K_{V})$ denote their standard neighbourhoods.
Figure 6. Left: Adding successive bypasses produces a sequence of (red) handlbodies $U_{i}$ that are homeomorphic but not weakly contact isotopic to $\nu(K_{U})$. Right: Adding the $1$-handles produces a bridge splitting $\widehat{H}=(\widehat{\Sigma},\widehat{U},\widehat{V})$; here $\widehat{U}$ is indicated in blue.
Begin with a smooth Heegaard splitting $\mathcal{H}=(\Sigma,U,V)$. Since the complement $X=M\setminus\big{(}\nu(K_{U})\cup\nu(K_{V})\big{)}$ is topologically a product, Theorem 2.6 implies that up to weak contact isotopy, $X$ decomposes as a sequence of bypasses $X_{1}\ast\dots\ast X_{f}$ attached to $\partial\nu(K_{U})$.111the triple $(\nu(K_{U}),\nu(K_{V})),X)$ is called a ”patty” in [HBH]. Attaching only the first $i$ bypasses in the sequence determines a new handlebody $U_{i}$ with convex boundary; letting $i$ range over the indices of all bypasses produces a sequence of smoothly isotopic Heegaard surfaces $\Sigma_{i}:=\partial U_{i}$. Each $\Sigma_{i}$ has a bypass arc $c_{i+1}$ that determines the bypass $X_{i+1}$. Isotopy of $c_{i}$ through bypass arcs preserves the weak contact isotopy class of $U_{i}$, so we may push each bypass arc off the faces of any $2$-handles already attached.
Definition 3.6.
Let $X_{1}\ast\dots\ast X_{f}$ be a bypass decomposition of $M\setminus\big{(}\nu(K_{U})\cup\nu(K_{V})\big{)}$ as above. Define $\widehat{U}$ to be the handlebody formed by adding only the bypass $1$-handles to $\nu(K_{U})$:
$\widehat{U}=\nu(K_{U})\cup\big{(}\bigcup_{i=1}^{f}h^{1}_{i}\big{)}.$
Then the bridge of $\mathcal{H}$ is the Heegaard splitting
^$\widehat{\mathcal{H}}(X_{1}\ast\dots\ast X_{f})=\big{(}\partial\widehat{U},\widehat{U},M\setminus\text{int}(\widehat{U})\big{)}.$
Attaching a contact $1$-handle preserves tightness, so it is immediate that $\widehat{U}$ is tight. The complementary $\widehat{V}$ is also tight, as turning the manifold upside down builds $\widehat{V}$ from $1$-handles attached to $\nu(K_{V})$. In fact, the splitting ^$\widehat{\mathcal{H}}(X_{1}\ast\dots\ast X_{f})=(\widehat{\Sigma},\widehat{U},\widehat{V})$ is convex, as co-core discs for each $1$-handle, together with meridional discs for the neighbourhoods of the skeleta, form a complete set of product discs.
Remark 3.7.
Although the term “bridge” is new, this construction appears in both [HBH] and [LV] as a technique to construct a convex Heegaard splitting from an arbitrary smooth Heegaard splitting.
When the original Heegaard surface $\Sigma$ is already convex in the splitting $\mathcal{H}=(\Sigma,U,V)$, one may consider the relative bridge ^$\widehat{\mathcal{H}}^{U}$ built from a bypass decomposition of $M\setminus\big{(}\nu(K_{U})\cup V\big{)}$, although this need not be convex. Similarly, ^$\widehat{\mathcal{H}}^{V}$ is constructed from a bypass decomposition of $M\setminus\big{(}U\cup\nu(K_{V})\big{)}$. The relative bridges will be used in Proposition 3.10.
Proposition 3.8.
Up to positive stabilisation, the bridge
^$\widehat{\mathcal{H}}(X_{1}\ast\dots\ast X_{f})=(\widehat{\Sigma},\widehat{U},\widehat{V})$
is independent of the choice of bypass attaching arcs within their isotopy classes.
Proof.
In order to construct a bridge splitting from a bypass decomposition, all the bypass attaching arcs must be isotoped off the faces of an $2$-handles already attached, but there is no canonical way to perform this isotopy. Although different isotopies preserve the weak contact isotopy type of each $U_{i}$, they may produce handlebodies which are inequivalent when the 2-handles are removed. We will show that an isotopy of the attaching 00-sphere for a $1$-handle on $\Sigma_{i}$ preserves the positive stabilisation class of the bridge $\widehat{\mathcal{H}}$.
Figure 7. Left: Before constructing the bridge of $\mathcal{H}$, bypass arcs must be isotoped off the faces of the $2$-handles. Centre: Isotoping $c$ across the face of a $2$-handle on $\Sigma_{i}$ preserves the weak contact isotopy class of $U_{i}$. Right: When $h_{i}^{2}$ is removed, the bypass arcs $c^{R}$ and $c^{L}$ are not isotopic on $\widehat{\Sigma}$.
Figure 8. After attaching a bypass along $c_{T}$, the arcs $c^{R}$ and $c^{L}$ again become isotopic.
Let $c^{R}$ and $c^{L}$ be two bypass attaching arcs for some $X_{j}$ ($j>i$) that are isotopic on $\Sigma_{i}$ via an isotopy that slides one endpoint across a face of some $2$-handle. Because the proof is local, we may assume without loss of generality that the associated bypass is the most recent one; that is, $j=i+1$. See Figure 7. The Heegaard surfaces $\widehat{\Sigma}^{R}$ and $\widehat{\Sigma}^{L}$ associated to attaching bypasses along $c^{R}$ and $c^{L}$, respectively, are not convexly isotopic.
Figure 9. Left: The bypass arcs $c^{R}$ and $c^{L}$ are isotopic on $\Sigma_{i}$. Centre: If $h_{i}^{2}$ is not attached, then $c^{R}$ and $c^{L}$ are not isotopic. Right: After attaching a bypass along $c_{T}$, the arcs $c^{R}$ and $c^{L}$ are again isotopic.
Consider the bypass arc $c_{T}$ shown on Figure 9. As the endpoints of $c^{L}$ and $c^{R}$ can be assumed to be arbitrarily close to $c_{i}$, we can also assume that $c_{T}$ is close to $c_{i}$. Thus no other handles are attached on top of $c_{T}$ and it persists as a bypass arc on $\widehat{\Sigma}^{L}$. We claim that there is a bypass half-disc attached along $c^{T}$ in $\widehat{V}^{L}$.
Observe that on $\Sigma_{i}$, the bypass arc $c_{T}$ is a rotation of $c_{i}$. As discussed in Section 2.3, there must exist a bypass disc $D_{T}$ attached along $c_{T}$. Moreover, by Lemma 2.8 $D_{T}$ may be assumed to lie in a neighbourhood of $X_{i}$ and the shaded area $R^{\prime}$ in the centre picture of Figure 9. It follows that the $1$-handle attached along $\partial c_{L}$ is disjoint from $D_{T}$, so $D_{T}$ persists in $\widehat{V}^{L}$.
Figure 9 shows that attaching the bypass along $c_{T}$ onto $\widehat{U}^{L}$ produces a handlebody weakly contact isotopic to $\widehat{U}^{R}$. Then by Theorems 2.5 and 3.4, the convex Heegaard decompositions ^$\widehat{\mathcal{H}}^{L}$ and ^$\widehat{\mathcal{H}}^{R}$ admit a common positve stabilisation. ∎
Proposition 3.8 implies that up to positive stabilisation, the bridge splitting ^$\widehat{\mathcal{H}}=(\widehat{\Sigma},\widehat{U},\widehat{V})$ depends only on $K_{U}$, $K_{V}$, and the bypass decomposition $X=X_{1}\ast\cdots\ast X_{f}$. As we show next, the bridge $\widehat{\mathcal{H}}$ is actually independent – again, up to positive stabilisation – of the bypass decomposition.
Proposition 3.9.
Given a smooth Heegaard splitting $\mathcal{H}=(\Sigma,U,V)$ let $K_{U}$ and $K_{V}$ be Legendrian skeletons for $U$ and $V$ and let $X=M\setminus\big{(}\nu(K_{U})\cup\nu(K_{V})\big{)}$. Let $X_{1}\ast\dots\ast X_{f}$ and $X^{\prime}_{1}\ast\dots\ast X^{\prime}_{f^{\prime}}$ be distinct bypass decompositions of $X$, and let ^=^$\widehat{\mathcal{H}}=\widehat{\mathcal{H}}(X_{1}\ast\dots\ast X_{f})$ and ^=^$\widehat{\mathcal{H}}^{\prime}=\widehat{\mathcal{H}}^{\prime}(X^{\prime}_{1}\ast\dots\ast X^{\prime}_{f})$ be the associated bridge Heegaard splittings formed by attaching only the $1$-handles of the respective bypasses to $\nu(K_{U})$. Then $\widehat{\mathcal{H}}$ and $\widehat{\mathcal{H}}^{\prime}$ admit a common positive stabilisation.
Proof.
According to Theorem 2.7, two bypass decompositions of $X$ are related by a finite sequence of isotopies, Trivial Insertion moves, and Far Commutation moves. Up to an isotopy that keeps the Heegaard surface convex, Far Commutation preserves the handlebodies $\widehat{U}$ and $\widehat{V}$. Thus we only need to show that inserting a trivial bypass preserves the positive stabilisation class of $\widehat{\mathcal{H}}$.
Fix a bypass decomposition of $X$. As above, let ^$\widehat{\mathcal{H}}=(\widehat{\Sigma},\widehat{U},\widehat{V})$ be the Heegaard splitting formed by attaching the associated $1$-handles to $\nu(K_{U})$. Now consider an alternative bypass decomposition that differs by the insertion of a trivial bypass $X_{T}=h^{1}_{T}\cup h^{2}_{T}$ attached along a trivial bypass arc $c_{T}$ somewhere in the sequence, and let ^$\widehat{\mathcal{H}}_{T}=(\widehat{\Sigma}_{T},\widehat{U}_{T},\widehat{V}_{T})$ denote the associated Heegaard splitting. Observe that the genus of ^$\widehat{\mathcal{H}}_{T}$ is larger than the genus of $\widehat{\mathcal{H}}$ because the additional $1$-handle $h_{T}^{1}$ is attached.
We claim that regardless of where the trivial bypass was inserted, $\widehat{\mathcal{H}}$ and ^$\widehat{\mathcal{H}}_{T}$ admit a common positive stabilisation. Recall that Proposition 3.8 allows us to isotope each attaching sphere along respective components of $\Gamma_{\Sigma_{i}}$ while preserving the positive stabilisation class of the bridge splitting. Using this flexibility, ensure that no handle attached after $h_{T}^{1}$ has an attaching sphere that intersects the disc supporting the trivial bypass. Not only can any attaching sphere be isotoped off this region, but we may also impose the stronger condition that subsequent attaching spheres are disjoint from the larger disc shown in Figure 10 that contains the bigon certifying $c_{T}$ as a trivial bypass arc. As a consequence, $c_{T}$ remains a trivial bypass arc on $\widehat{\Sigma}$; we will use this below. It follows that we may apply a sequence of Far Commutation moves to get a new bypass decomposition in which the bypass along $c_{T}$ is attached last.
Figure 10. Left: The purple arc is a trivial bypass attachment arc. Right: Attaching spheres for handles attached after this bypass may always be isotoped off the shaded gray disc.
Attaching $h^{1}_{T}$ along $\partial c_{T}$ to $\widehat{U}$ produces the handlebody $\widehat{U}_{T}$. As noted in the comment before Theorem 3.4, adding a bypass $1$-handle to a convex splitting is always a positive stabilisation, which proves the stated claim.
The following two results show that a convex Heegaard splitting $\mathcal{H}$ and its bridge $\widehat{\mathcal{H}}$ admit a common positive stabilisation.
Proposition 3.10.
Given a tight Heegaard splitting $\mathcal{H}=(\Sigma,U,V)$, let $K_{U}$, $K_{V}$ be Legendrian skeletons for $U$ and $V$, respectively. Let $X=X_{1}\ast\cdots\ast X_{f}$ be a bypass decomposition of the complement of $\nu(K_{V})$ in $V$, where the index is the order in which the bypass slices are attached to $U$. Let ^$\widehat{\mathcal{H}}^{U}$ denote the relative bridge splitting with handlebodies $\widehat{U}=U\cup\big{(}\bigcup_{i=1}^{f}h^{1}_{i}\big{)}$ and $\widehat{V}=M\setminus\text{int}(\widehat{U})$. Then the refinements of the Heegaard splittings ^$\widehat{\mathcal{H}}^{U}$ and $\mathcal{H}$ admit a common positive stabilisation.
The relative bridge splitting ^$\widehat{\mathcal{H}}^{U}$ is tight by construction, so one may flip this Heegaard splitting upside down and apply Proposition 3.10 again to get the following consequence:
Corollary 3.11.
Given a tight Heegaard splitting $\mathcal{H}=(\Sigma,U,V)$ let $K_{U}$, $K_{V}$ be Legendrian skeletons for $U$ and $V$ and let $X=X_{1}\ast\cdots\ast X_{f}$ be a bypass decomposition of the complement of $M\setminus\big{(}\nu(K_{U})\cup\nu(K_{V})\big{)}$. Then the bridge splitting $\widehat{\mathcal{H}}$ and the refinement $\widetilde{\mathcal{H}}$ admit a common positive stabilisation.
The proof of Proposition 3.10 is essentially the same argument as the proof of [LV, Theorem 6.1.]. Since it was not stated in full generality there, we indicate how it follows from the techniques of [LV].
Proof of Proposition 3.10..
Define the sequence of intermediate Heegaard splittings ^$\widehat{\mathcal{H}}_{j}=(\widehat{\Sigma}_{j},\widehat{U}_{j},\widehat{V}_{j})$ built by attaching only the first $j$ of the $1$-handles:
$\widehat{U}_{j}=U\cup\big{(}\bigcup_{i=1}^{j}h^{1}_{i}\big{)},\ \ \widehat{V}_{j}=M\setminus\text{int}(\widehat{U}_{j}),\ \ \widehat{\Sigma}_{j}=\partial\widehat{U}_{j}.$
We will show that the refinements of successive Heegaard splittings in this sequence admit common positive stabilisations.
Beginning with the case $j=1$, construct the refinement ~$\widetilde{\mathcal{H}}=(\widetilde{\Sigma},\widetilde{U},\widetilde{V})$ of $\mathcal{H}$.
Recall that the refinement is constructed by drilling tunnels through meridional discs for the handlebodies. We may choose these discs to be disjoint from the bypass $X_{1}=h^{1}_{1}\cup h^{2}_{1}$, so adding the $1$-handle $h_{1}^{1}$ to $U$ commutes with this drilling. Thus attaching $h^{1}_{1}$ to the refined $\widetilde{U}$ produces the same Heegaard splitting as taking the refinement of ^$\widehat{\mathcal{H}}_{1}$. As noted in Section 3.1, attaching $h_{1}^{1}$ in to $\widetilde{U}$ in the convex refinement is a positive stabilisation, which completes the base case.
Now proceed by induction on $j$. For the inductive step, note that if ^$\widehat{\mathcal{H}}_{j}$ is tight, then ^$\widehat{\mathcal{H}}_{j+1}$ is also tight because $\widehat{V}_{j+1}$ is a proper subset of the tight $\widehat{V}_{j}$, while $\widehat{U}_{j+1}$ is built from the tight $\widehat{U}_{j}$ by adding a contact $1$-handle. The argument above shows that refining and then adding a $1$-handle produces the same splitting as first adding the $1$-handle and then refining, which preserves the positive stabilisation class. This establishes the inductive step. ∎
3.3. Proof of the Giroux Correspondence
In order to prove the Giroux Correspondence, we consider a pair of convex Heegaard splittings of a fixed contact manifold. We will modify these Heegaard splittings until they coincide, and each step in this evolution will preserve the positive stabilisation class of the Heegaard splittings.
First, we repeat an argument from the proof of Theorem 6.1 in [LV] that lets us restrict to Heegaard splittings that are smoothly isotopic:
Proposition 3.12.
Any two convex Heegaard splittings of $(M,\xi)$ have positive stabilisations that are smoothly isotopic.
Proof.
In the topological setting, the Reidemeister-Singer Theorem asserts that any two Heegaard splittings of a fixed manifold become isotopic after sufficiently many stabilisations. As there are no restrictions on the type of stabilisation that ensure this outcome, we are free to choose positive stabilisation of our convex Heegaard splittings. ∎
Next, we can assume that the skeletons of the Heegaard splittings agree:
Proposition 3.13.
Let $\mathcal{H}=(\Sigma,U,V)$ and $\mathcal{H}^{\prime}=(\Sigma^{\prime},U^{\prime},V^{\prime})$ be smoothly isotopic Heegaard splittings of $(M,\xi)$ with convex Heegaard surfaces. Then there exist Legendrian skeletons $K_{U}$ and $K_{V}$ and an isotopy $\Psi_{t}$ of $\mathcal{H}^{\prime}$ keeping $\Psi_{t}(\Sigma^{\prime})$ convex such that $K_{U}$ is the skeleton of both $U$ and $\Psi_{1}(U^{\prime})$ and $K_{V}$ is the skeleton of both $V$ and $\Psi_{1}(V)$.
Proof.
Choose a Legendrian bouquet $K_{*}$ as a skeleton for each handlebody, where $*\in\{U,U^{\prime},V,V^{\prime}\}$. Since $K_{{U^{\prime}}}\cup K_{{V^{\prime}}}$ is topologically isotopic to $K_{{U}}\cup K_{{V}}$, some stabilisations of these bouquets are Legendrian isotopic. As we may stabilise each $K_{*}$ while preserving it as a skeleton of the respective handlebody, suppose now that the skeletons $K_{{U^{\prime}}}\cup K_{{V^{\prime}}}$ are Legendrian isotopic to $K_{{U}}\cup K_{{V}}$.
Extend this Legendrian isotopy to an ambient contact isotopy of the manifold, $\Psi_{t}$. By construction, for all $t$ the bouquet $\Psi_{t}(K_{{U^{\prime}}})$ is a skeleton for the handlebody $\Psi_{t}({{U^{\prime}}})$ with convex boundary $\Psi_{t}(\Sigma)$, and the analogous statement holds for $\Psi_{t}(K_{{V^{\prime}}})$. Since $\Psi_{t}$ is an extension of the Legendrian isotopy taking $K_{{U^{\prime}}}\cup K_{{V^{\prime}}}$ to $K_{{U}}\cup K_{{V}}$, we also have that $\Psi_{1}(K_{{U^{\prime}}})=K_{{U}}$ is a skeleton for $U$ and $\Psi_{1}(K_{{V^{\prime}}})=K_{{V}}$ is a skeleton for $V$, as desired. ∎
Finally, we are ready to prove the Giroux Correspondence for an arbitrary contact $3$-manifold.
Proof of Theorem 3.3.
Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be two convex Heegaard splittings of $(M,\xi)$. By Proposition 3.12 we may assume that these Heegaard decompositions are smoothly isotopic. Then Proposition 3.13 implies that after performing an isotopy keeping $\Sigma^{\prime}$ convex, we may also assume there exist a Legendrian bouquet $K_{U}$ that is a skeleton of both $U$ and $U^{\prime}$ and a Legendrian bouquet $K_{V}$ that is a skeleton of both $V$ and $V^{\prime}$.
Choose bypass decompositions of the complement of the skeletons in $U$ and $V$:
${U\setminus\nu(K_{U})}=X_{1}\ast\cdots\ast X_{l}\text{ \ \ \ and \ \ \ }{V\setminus\nu(K_{V})}=X_{l+1}\ast\cdots\ast X_{k}.$
Concatenating these produces a bypass decomposition of $X=M\setminus\big{(}\nu(K_{U})\cup\nu(K_{V})\big{)}$. As usual, let $\widehat{U}$ denote the handlebody formed by attaching all of the $1$-handles to $\nu(K_{U})$. Setting ^$\widehat{\mathcal{H}}=(\partial\widehat{U},\widehat{U},M\setminus\text{int}\widehat{U})$, Corollary 3.11 implies that the refinement $\widetilde{\mathcal{H}}$ and the bridge $\widehat{\mathcal{H}}$ admit a common positive stabilisation. To complete the argument, we recall Remark 3.5: since $\mathcal{H}$ is convex, any refinement of $\widetilde{\mathcal{H}}$ is a positive stabilisation of $\mathcal{H}$ itself. Thus $\mathcal{H}$ and $\widehat{\mathcal{H}}$ admit a common positive stabilisation.
Now repeat this process using the alternative Heegaard splitting $\mathcal{H}^{\prime}$:
${U^{\prime}\setminus\nu(K_{U})}=X^{\prime}_{1}\ast\cdots\ast X^{\prime}_{l^{\prime}}\text{ \ \ \ and \ \ \ }V^{\prime}\setminus\nu(K_{V})=X^{\prime}_{l^{\prime}+1}\ast\cdots\ast X^{\prime}_{k^{\prime}}.$
Attaching all – and only – the $1$-handles to $\nu(K_{U})$ gives a new bridge splitting $\widehat{\mathcal{H}}^{\prime}$, and a similar argument shows that $\mathcal{H}^{\prime}$ and $\widehat{\mathcal{H}}^{\prime}$ admit a common positive stabilisation.
However, $\widehat{\mathcal{H}}$ and $\widehat{\mathcal{H}}^{\prime}$ are built from distinct bypass decompositions of $M\setminus\big{(}\nu(K_{U})\cup\nu(K_{V})\big{)}$, so Proposition 3.9 implies that they also admit a common positive stabilisation, completing the proof. ∎